# A Lensing-Band Approach to Spacetime Constraints

### 11 Dec 2023 v1

@@ Acknowledgements.
We thank C. Bambi, A. Chael, A. Eichhorn, D. Gates, R. Gold, P. Kocherlakota, A. Lupsasca, D. Mayerson, F. Pretorius, F. Vincent, and G. Wong for their comments. A.H. thanks the Princeton Gravity Initiative for their hospitality during substantial parts of this work. A.H. acknowledges support from the PRIME programme of the German Academic Exchange Service (DAAD) with funds from the German Federal Ministry of Education and Research (BMBF). A.H. also acknowledges support by the Deutsche Forschungsgemeinschaft (DFG) under Grant No 406116891 within the Research Training Group RTG 2522/1. A.C.-A. acknowledges support from Will and Kacie Snellings and the Simons Foundation. Some of the simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University. Map colors were based on www.ColorBrewer.org, by Cynthia A. Brewer, Penn State. In preparing this manuscript we made use of the python packages numpy [122], scipy [123], and matplotlib [124].
@@ Appendix A Free-floating bisection algorithm
Here, we provide the details of an algorithm to solve the following numerical problem: We assume a 2D vector space $\mathbb{R}^{2}$ with a suitable norm ||$\cdot$|| ⋅ |, and a compact subregion $\mathcal{R}\subset\mathbb{R}^{2}$ defined by a binary condition $\mathcal{C}$ such that $\mathcal{C}(\vec{x})=\text{TRUE}$ if $\vec{x}\in\mathcal{R}$ and $\mathcal{C}(\vec{x})=\text{FALSE}$ otherwise. The goal of the algorithm is to numerically approximate one closed piece of the boundary $\partial\mathcal{R}$ of this region. If the region is topologically nontrivial, there may be several such closed pieces which will then have to be obtained individually [^12] For the present use-case of a lensing-band region, there are two distinct closed pieces of boundary—one inner and one outer boundary of the lensing band..
@@
    Determine an initial section of the boundary: First, we need to obtain two initial points $\vec{x}_{1},\,\vec{x}_{2}\in\mathbb{R}^{2}$ which approximate two distinct, but sufficiently close, boundary points. With the term “distinct, but sufficiently close”, we refer to $0<|\vec{x}_{2}-\vec{x}_{1}|<\epsilon$ such that $\epsilon$ is smaller than the desired (user-defined) initial numerical precision. This step either requires some prior knowledge about where (at least a piece of) the boundary is located, or needs to be iterated by drawing random pairs of $\vec{x}_{1},\,\vec{x}_{2}\in\mathbb{R}^{2}$ until an initial boundary section has been successfully identified. We detail in the main text how this initial step can be robustly implemented for the specific application to lensing bands.
@@
    Take a free-floating bisection step (iterated): Given the initial section of the boundary (defined by $\vec{x}_{1}$ and $\vec{x}_{2}$), the precision $p$, and the initial inner and outer guard-rail points $\vec{i}$ and $\vec{o}$, we can now perform the key iterative step of the algorithm. This step is further subdivided into the following tasks.
@@
      First, we normalize the tangent by the precision $p$.
@@
      Second, we move forward either $\vec{i}$ or $\vec{o}$ along the tangent $\vec{t}$. In the first iteration this choice is arbitrary but in each subsequent step we determine whether to move forward $\vec{i}$ and $\vec{o}$ depending on whether the scalar product $\vec{t}\cdot\vec{n}$ is positive or negative. This ensures that each bisection step will (approximately) follow the curvature of the boundary.
@@
      Third, we perform a standard boundary bisection between the updated $\vec{i}^{\prime}$ and $\vec{o}^{\prime}$. If this bisection converges (up to precision $p$) to either of the guardrail points, we can no longer be sure that the boundary remains between the guardrail points. If the latter occurs, we increase the precision $p^{\prime}=a_{\text{tighten}}\,p$ by a factor $a_{\text{tighten}}<1$ and repeat the above three tasks. This increase in the precision eventually ensures that the boundary remains between the guardrail points.
@@ Appendix B The KRZ expansion of circular spacetimes
In this Appendix we briefly review the expansion of circular spacetimes provided in Ref. [113]. For specific details, we refer the reader to Refs. [112, 113]. This particular parametrization uses both polynomial and continued-fraction expansions to represent the metric in (7) by five free functions $B$, $K$, $N$, $W$, and $\Sigma$, i.e.,
@@
which is compatible with a representation of Kerr spacetime in Boyer–Lindquist coordinates, cf. $\sigma(r,\theta)$ below Eq. (7). In the latter case, $A_{\text{KRZ}}$ matches the spin parameter of Kerr spacetime. More generally, the parameter $A_{\text{KRZ}}$ is related to the asymptotic choice of coordinates, i.e., to the focal parameter of the oblate spheroidal coordinates in which the asymptotically flat limit of the spacetime is represented–see Appendix C for further discussion.
@@
We summarize the explicit matching to the Kerr solution in Tab. 1 and in App. B. Note that the matching to the Kerr metric presented in Ref. [113] needs additional terms that are not shown explicitly in their paper [125]. The last term in [113, Eq. (A6)] should read
@@
The coefficients $k_{i0}$ play a special role because they are not suppressed by $r_{0}/r$ and thus modify the oblate spheroidal coordinates to which the Boyer–Lindquist coordinated converge in the flat asymptotic limit, as shown in Appendix C. We choose to fix all the $k_{i0}$ as in Ref. [113], i.e.,
@@
such that the parameter $A_{\text{KRZ}}$ corresponds to the focal parameter of the asymptotically flat oblate spheroidal coordinates (cf. Appendix C). The coefficients $\epsilon_{i}$, $b_{i0}$, $\omega_{i0}$, and $k_{i0}$ contribute at order $\mathcal{O}(r_{0}/r)$, while $a_{i0}$ contributes at order $\mathcal{O}(r_{0}^{2}/r^{2})$, cf. also Table 1.
@@
Thus, with regards to the leasing asymptotics, the physical parameters of the metric in this parametrization can be identified by trading $(\delta\epsilon_{0},\,\delta\omega_{00})$ for $(M,\,J)$ (using Eqs (44)) and $(\delta a_{00},\,\delta b_{00})$ for $(\beta,\,\gamma)$ (using Eqs. (45)), to get
@@
If one assumes uniqueness and Solar System constraints on the PPN parameters, $\gamma$ and $\beta$ are required to be close to one [7]. Higher-order coefficients correspond to higher-order corrections to the behaviour at asymptotic infinity. The mapping to the PPN metric of a point particle is shown explicitly in Appendix C.
@@
Table 1: The $51$ parameters $(\epsilon_{i},\,a_{ij},\,b_{ij},\,w_{ij},\,k_{ij})$, up to $\mathcal{O}(x^{3}y^{2})$, ( $i\leqslant 2$ and $j\leqslant 3$), required in the KRZ expansion to represent the Kerr spacetime exactly. The KRZ parameters $\epsilon_{i}$ and those with $j=0$ are chosen to obtain a specific form of asymptotics. In contrast, the KRZ parameters with $j\geqslant 1$ are part of a continued-fraction expansion. Once any of the continued-fraction parameters vanishes, all corresponding higher-order parameters do not contribute. The KRZ parameters with odd $i$ break reflection symmetry about the equatorial plane, and thus all vanish for Kerr spacetime. We do not consider such deviations in the present work. We list the respective values reproducing the Kerr spacetime, where $r_{0}=m+\sqrt{m^{2}-a^{2}}$. We also indicate the order in the $1/r$ expansion around $r=\infty$ at which the respective deviation parameter affects the asymptotic behaviour of the spacetime by $\mathcal{O}(r^{-n})$. All the coefficients for which $n=0$ affect the asymptotic definition of oblate spheroidal coordinates, cf. Appendix C. Finally, we highlight in bold the $5$ deviation parameters which are investigated in the present work. Some of the leading and subleading continued-fraction parameters have also been investigated in Refs. [48, 47].
@@
As a result, we obtain a spacetime with asymptotic mass and angular momentum $(M,\,J)$ and $5$ deviation parameters $(r_{0},\,s,\,\beta,\,\gamma,\,\delta a_{01})$. When $r_{0}=M+\sqrt{M^{2}-J^{2}/M^{2}}$, $a=J/M$, $\beta=\gamma=1$, and $\delta a_{01}=0$, the spacetime reduces to Kerr spacetime in Boyer–Lindquist coordinates. More generally, its metric elements are compactly written as
@@
It is $A_{\text{KRZ}}$, not $A=J/M$ or $a$ which appears in these shorthands. In addition, we use additional deviation shorthands, i.e.,
@@ C.1 The far-field metric in GR in quasi-isotropic coordinates
Other uniqueness theorems extend the importance of the above asymptotic behavior to the vacuum exterior of generic spacetimes in two ways. First, in spherical symmetry, i.e., for $J\rightarrow 0$, Birkhoff’s theorem guarantees that the same asymptotic behavior applies to the vacuum exterior of any source [1]. Second, the leading-order (i.e., up to $\mathcal{O}(1/r^{2})$) behavior also applies to the vacuum exterior of any stationary and axisymmetric source [3].
@@ Appendix D Theoretical constraints on the KRZ parameters
    We thus need to make sure that $r=r_{0}$ remains the largest root of Eq. (58).

### 05 Mar 2024 v2

@@ Acknowledgements.
We thank C. Bambi, A. Chael, A. Eichhorn, D. Gates, R. Gold, P. Kocherlakota, A. Lupsasca, D. Mayerson, F. Pretorius, F. Vincent, and G. Wong for their comments. We also thank an anonymous referee for their constructive comments. A.H. thanks the Princeton Gravity Initiative for their hospitality during substantial parts of this work. A.H. acknowledges support from the PRIME programme of the German Academic Exchange Service (DAAD) with funds from the German Federal Ministry of Education and Research (BMBF). A.H. also acknowledges support by the Deutsche Forschungsgemeinschaft (DFG) under Grant No 406116891 within the Research Training Group RTG 2522/1. A.C.-A. acknowledges support from Will and Kacie Snellings and the Simons Foundation. Some of the simulations presented in this work were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University. Map colors were based on www.ColorBrewer.org, by Cynthia A. Brewer, Penn State. In preparing this manuscript we made use of the python packages numpy [126], scipy [127], and matplotlib [128].
@@ Appendix A Free-floating bisection algorithm
Here, we provide the details of an algorithm to solve the following numerical problem: we assume a 2D vector space $\mathbb{R}^{2}$ with a suitable norm ||$\cdot$|| ⋅ |, and a compact subregion $\mathcal{R}\subset\mathbb{R}^{2}$ defined by a binary condition $\mathcal{C}$ such that $\mathcal{C}(\vec{x})=\text{TRUE}$ if $\vec{x}\in\mathcal{R}$ and $\mathcal{C}(\vec{x})=\text{FALSE}$ otherwise. The goal of the algorithm is to numerically approximate one closed piece of the boundary $\partial\mathcal{R}$ of this region. If the region is topologically nontrivial, then there may be several such closed pieces which will then have to be obtained individually [^12] For the present use-case of a lensing-band region, there are two distinct closed pieces of boundary—one inner and one outer boundary of the lensing band..
@@
    Determine an initial section of the boundary: First, we need to obtain two initial points $\vec{x}_{1},\,\vec{x}_{2}\in\mathbb{R}^{2}$ which approximate two distinct, but sufficiently close, boundary points. With the phrase “distinct, but sufficiently close”, we refer to $0<|\vec{x}_{2}-\vec{x}_{1}|<\epsilon$ such that $\epsilon$ is smaller than the desired (user-defined) initial numerical precision. This step either requires some prior knowledge about where (at least a piece of) the boundary is located, or needs to be iterated by drawing random pairs of $\vec{x}_{1},\,\vec{x}_{2}\in\mathbb{R}^{2}$ until an initial boundary section has been successfully identified. We detail in the main text how this initial step can be robustly implemented for the specific application to lensing bands.
@@
    Take a free-floating bisection step (iterated): Given the initial section of the boundary (defined by $\vec{x}_{1}$ and $\vec{x}_{2}$), the precision $p$, and the initial inner and outer guard-rail points $\vec{i}$ and $\vec{o}$, we can now perform the key iterative step of the algorithm. This step is further subdivided into the following tasks:
@@
      We normalize the tangent by the precision $p$.
@@
      We move forward either $\vec{i}$ or $\vec{o}$ along the tangent $\vec{t}$. In the first iteration this choice is arbitrary but in each subsequent step we determine whether to move forward $\vec{i}$ and $\vec{o}$ depending on whether the scalar product $\vec{t}\cdot\vec{n}$ is positive or negative. This ensures that each bisection step will (approximately) follow the curvature of the boundary.
@@
      We perform a standard boundary bisection between the updated $\vec{i}^{\prime}$ and $\vec{o}^{\prime}$. If this bisection converges (up to precision $p$) to either of the guardrail points, then we can no longer be sure that the boundary remains between the guardrail points. If the latter occurs, then we increase the precision $p^{\prime}=a_{\text{tighten}}\,p$ by a factor of $a_{\text{tighten}}<1$ and repeat the above three tasks. This increase in the precision eventually ensures that the boundary remains between the guardrail points.
@@ Appendix B The KRZ expansion of circular spacetimes
In this Appendix we briefly review the expansion of circular spacetimes provided in Ref. [116]. For specific details, we refer the reader to Refs. [115, 116]. This particular parametrization uses both polynomial and continued-fraction expansions to represent the metric in Eq. 7 by five free functions $B$, $K$, $N$, $W$, and $\Sigma$, i.e.,
@@
which is compatible with a representation of Kerr spacetime in Boyer–Lindquist coordinates, cf. $\sigma(r,\theta)$ below Eq. 7. In the latter case, $A_{\text{KRZ}}$ matches the spin parameter of Kerr spacetime. More generally, the parameter $A_{\text{KRZ}}$ is related to the asymptotic choice of coordinates, i.e., to the focal parameter of the oblate spheroidal coordinates in which the asymptotically flat limit of the spacetime is represented–see App. C for further discussion.
@@
We summarize the explicit matching to the Kerr solution in Tab. 1 and in App. B. Note that the matching to the Kerr metric presented in Ref. [116] needs additional terms that are not shown explicitly in their paper [129]. The last term in [116, Eq. (A6)] should read
@@
The coefficients $k_{i0}$ play a special role because they are not suppressed by $r_{0}/r$ and thus modify the oblate spheroidal coordinates to which the Boyer–Lindquist coordinated converge in the flat asymptotic limit, as shown in App. C. We choose to fix all the $k_{i0}$ as in Ref. [116], i.e.,
@@
such that the parameter $A_{\text{KRZ}}$ corresponds to the focal parameter of the asymptotically flat oblate spheroidal coordinates (cf. App. C). The coefficients $\epsilon_{i}$, $b_{i0}$, $\omega_{i0}$, and $k_{i0}$ contribute at order $\mathcal{O}(r_{0}/r)$, while $a_{i0}$ contributes at order $\mathcal{O}(r_{0}^{2}/r^{2})$, cf. also  Tab. 1.
@@
Thus, with regards to the leasing asymptotics, the physical parameters of the metric in this parametrization can be identified by trading $(\delta\epsilon_{0},\,\delta\omega_{00})$ for $(M,\,J)$ (using Eqs. 44) and $(\delta a_{00},\,\delta b_{00})$ for $(\beta,\,\gamma)$ [using Eqs. 45], to get
@@
If one assumes uniqueness and Solar System constraints on the PPN parameters, $\gamma$ and $\beta$ are required to be close to one [8]. Higher-order coefficients correspond to higher-order corrections to the behaviour at asymptotic infinity. The mapping to the PPN metric of a point particle is shown explicitly in App. C.
@@
Table 1: The $51$ parameters $(\epsilon_{i},\,a_{ij},\,b_{ij},\,w_{ij},\,k_{ij})$, up to $\mathcal{O}(x^{3}y^{2})$, ( $i\leqslant 2$ and $j\leqslant 3$), required in the KRZ expansion to represent the Kerr spacetime exactly. The KRZ parameters $\epsilon_{i}$ and those with $j=0$ are chosen to obtain a specific form of asymptotics. In contrast, the KRZ parameters with $j\geqslant 1$ are part of a continued-fraction expansion. Once any of the continued-fraction parameters vanishes, all corresponding higher-order parameters do not contribute. The KRZ parameters with odd $i$ break reflection symmetry about the equatorial plane, and thus all vanish for Kerr spacetime. We do not consider such deviations in the present work. We list the respective values reproducing the Kerr spacetime, where $r_{0}=m+\sqrt{m^{2}-a^{2}}$. We also indicate the order in the $1/r$ expansion around $r=\infty$ at which the respective deviation parameter affects the asymptotic behaviour of the spacetime by $\mathcal{O}(r^{-n})$. All the coefficients for which $n=0$ affect the asymptotic definition of oblate spheroidal coordinates, cf. App. C. Finally, we highlight in bold the five deviation parameters which are investigated in the present work. Some of the leading and subleading continued-fraction parameters have also been investigated in Refs. [49, 48].
@@
As a result, we obtain a spacetime with asymptotic mass and angular momentum $(M,\,J)$ and five deviation parameters $(r_{0},\,s,\,\beta,\,\gamma,\,\delta a_{01})$. When $r_{0}=M+\sqrt{M^{2}-J^{2}/M^{2}}$, $a=J/M$, $\beta=\gamma=1$, and $\delta a_{01}=0$, the spacetime reduces to Kerr spacetime in Boyer–Lindquist coordinates. More generally, its metric elements are compactly written as
@@
It is $A_{\text{KRZ}}$, not $A=J/M$ or $a$, which appears in shorthand. In addition, we use additional deviation shorthand, i.e.,
@@ C.1 The far-field metric in GR in quasi-isotropic coordinates
Other uniqueness theorems extend the importance of the above asymptotic behavior to the vacuum exterior of generic spacetimes in two ways. First, in spherical symmetry, i.e., for $J\rightarrow 0$, the Jebsen–Birkhoff theorem [1, 2] guarantees that the same asymptotic behavior applies to the vacuum exterior of any source [2]. Second, the leading-order (i.e., up to $\mathcal{O}(1/r^{2})$) behavior also applies to the vacuum exterior of any stationary and axisymmetric source [4].
@@ Appendix D Theoretical constraints on the KRZ parameters
    We thus need to make sure that $r=r_{0}$ remains the largest root of Eq. 58.